I have 7/6x - 3/2 = 6/9. How would you go about finding the answer?
_________________
Robert
There are people I know who won't hurt me. I call them corpses.
- Randy K. Milholland

Answer by checkley71(8403)

7/6x-3/2=6/9 or (14-18x)/12x=6/9 or 6*12x=9(14-18x) or 72x=126-162x or
234x=126 or x=126/234 or x=7/13